Here is the problem:
Quality Control:
Suppose that 10 computer chips are drawn from a box containing 12 good chips and 4 defective chips. What is the probability that
a. exactly 4 of the chips are defective
b. all 10 of the chips are good
c. exactly 8 of the chips are good
the answers are:
a. 924/8008
b. 66/8008
c. 2970/8008\
I tried using the nCr formula. I am sure this is the right one(not really sure or I would not be writing this) only i can't figure out which numbers to put where.
this is what I did:
working on a.
10!/4!(10-4)! = came up with 210
thanks in advance for any help.
Heidi
ps where do i find the thank you form. can't find it.

Answer by stanbon(48568)

Suppose that 10 computer chips are drawn from a box containing 12 good chips and 4 defective chips.
-------
These are binomial problems with p = 12/16 = 3/4, q = 1/4, n=10
What is the probability that
a. P(exactly 4 of the chips are defective) = 10C4(3/4)^6(1/4)^4
-------------------------------------------------
b. P(all 10 of the chips are good) = 10C10(3/4)^10(1/4)^0
------------------------------------------------
c. P(exactly 8 of the chips are good) = 10C8(3/4)^8(1/4)^8
-------------------
Comment: None of my answers agree with your posted answers.
-------------------
Cheers,
Stan H.
the answers are:
a. 924/8008
b. 66/8008
c. 2970/8008\